OpenStudy (anonymous):

Let R be the region of the plane bounded by the curves y = 2x and 6x − y = 8. The region R is rotated around the y-axis. Use the washer method to find an integral whose value is the volume of the solid that you get.

OpenStudy (anonymous):

Also assume bounded by x-axis?

OpenStudy (anonymous):

yes i think...please help

OpenStudy (anonymous):

I'm trying to remember difference between washer method, shell method, disk method, but either way, I think I would integrate along the y-axis using the area of the washer as \[\frac{y+8}{6} \space -\frac{y}{2}\]

OpenStudy (anonymous):

integrate that with a dy from 0 to 4. I think that ought to do the trick.

OpenStudy (anonymous):

Oh, sorry, I forgot the πr^2 business...

OpenStudy (anonymous):

why is it y+8/6 - y/2

OpenStudy (anonymous):

Just a sec, forgot to square the radii...

OpenStudy (anonymous):

I solved the equations for x instead of y, so I could integrate along the y-axis using dy instead of dx.

OpenStudy (anonymous):

I think that's what makes the washer shape instead of the shells shape, but I sometimes get those terms mixed up.

OpenStudy (anonymous):

Ok, sorry, the area of each washer is \[\left(\frac{y+8}{6} \right)^2 - \left(\frac{y}{2} \right)^2\]

OpenStudy (anonymous):

It helps if you graph the lines to see which is the outer radius and which is the inner radius. (If you haven't already)

OpenStudy (anonymous):

so can you please draw all the stuff so i know where to start

OpenStudy (anonymous):

Sure, I'll have to do it quick 'cause I'm about to leave to go to dinner.

OpenStudy (anonymous):

but i need all the work too...

OpenStudy (anonymous):



OpenStudy (anonymous):

You're going to have to do some of this work on your own . . .

OpenStudy (anonymous):

aw :(

OpenStudy (anonymous):

i can't

OpenStudy (anonymous):